3. Find the inverse of the function and state the domain of the inverse. $(x) = 4x Show that f(f(x) = x -6
Answer #1

# solution in the ginen function :- Jex) = 4x + (x-6 toy x. f(x) - 6 fox) = 4x x (f(x)-4)= 6 fexe) a 6t() feu)-4 64 f(x) = 6x